# Yuehao Bai

### v7

@@ 4.1 Efficiency Bound
Given Theorem 4.1 we call $\mathbb{V}_{\ast}=\operatorname{Var}[\psi^{\ast}(X_{i},A_{i},R_{i},\theta_{0})]$ the efficiency bound for $\theta_{0}$, since our result shows that this is the lowest asymptotic variance attainable by any regular estimator under our assumptions. Indeed, it follows from Anderson’s lemma (Lemma 8.5 in van_der_vaart1998asymptotic) that the asymptotic loss of any regular estimator is bounded below by the loss under $N(0,\mathbb{V}_{\ast})$ for any “bowl-shaped” loss function (including, in particular, square loss). We note that our assumptions on the assignment mechanism preclude us from immediately appealing to standard semi-parametric convolution theorems (see, for instance, Theorem 25.20 in van_der_vaart1989asymptotic). Instead, we proceed by justifying an application of Theorem 3.1 in armstrong2022asymptotic combined with the convolution Theorem 3.11.2 in van_der_vaart1996weak to each $d_{\theta}$-dimensional parametric submodel separately, and then arguing that the supremum over all such submodels is attained by $\operatorname{Var}[\psi^{\ast}]$. A key observation is that in order to apply Theorem 3.1 in armstrong2022asymptotic to argue that the likelihood ratio process is locally asymptotically normal, the conditional information needs to settle down in the limit, which is guaranteed as long as Assumption 4.1 is satisfied.
@@
Following similar arguments as those in Remark 3.2, we can deduce that our efficiency bound agrees with well-known bounds for common parameters (like those presented in Examples 2.12.3) in the setting of i.i.d. assignment. For example, we have noted in the case of the ATE (Example 2.1) that (17) matches the efficiency bound under i.i.d. assignment derived in hahn1998role. See rafi2023efficient and armstrong2022asymptotic for related results in the context of stratified and response-adaptive experiments. Straightforward calculation also implies that, for the quantile treatment effect (Example 2.2), the efficiency bound is given by
@@
which matches the efficiency bound under i.i.d. assignment derived in firpo2007efficient when the propensity score is set to $\eta$.
@@
The efficiency bound in Theorem 4.1 is attained by finely stratified experiments as in Theorem 3.1 if no additional covariates are available for estimation beyond the set of covariates $X_{i}$ used in the design. In practice, researchers may consider adjusting for additional baseline covariates in order to improve efficiency. Suppose additional covariates $W^{(n)}$ are available and Assumption 3.1 is modified such that
@@
Here, we comment on how Theorem 4.1 relates to prior efficiency bounds in experiments with general assignment mechanisms. For the case of estimating the ATE, armstrong2022asymptotic derives an efficiency bound over a very large class of assignment mechanisms, including even response-adaptive designs, and shows that the bound is attained when units are assigned to treatment (control) with conditional probability proportional to the conditional standard deviation of the potential outcome under treatment (control). This type of assignment is sometimes referred to as the Neyman allocation. On the other hand, our results show that this bound may be quite loose whenever the assignment proportions are restricted to be anything not equal to the Neyman allocation, which is, of course, unknown. For example, the bound is not informative about what can be achieved if the assignment proportions were set to one half regardless of whether or not the conditional outcome variances across treatment and control are equal. Such settings frequently arise in practice due to logistical constraints or the absence of pilot data with which to estimate conditional variances of potential outcomes under treatment and control. Furthermore, as argued in cai2022performance, even if pilot data is available, these quantities may be estimated so poorly that exogenously constraining the assignment proportions to one half leads to more efficient estimates of the ATE in practice. Motivated by such concerns, rafi2023efficient derives an efficiency bound for the ATE over the class of “coarsely-stratified” assignment mechanisms studied in bugni2019inference, where the stratum-level assignment proportions are restricted a priori by the experimenter. This framework, however, rules out finely stratified designs. Finally, we once again emphasize that our analysis, unlike these other papers, applies to a general class of treatment effect parameters, including the ATE as a special case.
@@
If we additionally impose that $\eta(X_{i})$ takes on a finite set of values $\{\eta_{1},\dots,\eta_{S}\}$, then this bound could be achieved by separately implementing a finely stratified experiment over each set $\{i:\eta(X_{i})=\eta_{s}\}$ for $1\leq s\leq S$. In other words, separately within each stratum defined by the units for which $\eta(X_{i})=\eta_{s}$, employ the assignment mechanism described in Assumptions 3.13.2 with $\ell/k=\eta_{s}$. For more general functions $\eta(\cdot)$, we conjecture one could employ the local randomization procedure in cytrynbaum2023designing.
@@ 4.2 The Necessity of “Fast-Balancing”
In this subsection, we provide conditions under which the fast-balancing condition described in (14) is a necessary condition for efficient estimation of $\theta_{0}$ “by design.” As a supplement, we also provide necessary and sufficient conditions for an asymptotically linear estimator to be regular (in the sense of (S.16) in Appendix A.2) for a large class of treatment assignment mechanisms. Concretely, given an assignment mechanism, suppose $\tilde{\theta}_{n}$ is an asymptotically linear estimator for $\theta_{0}$ in the sense that
@@
where $E[\psi(X_{i},A_{i},R_{i},\theta_{0})]=0$ and $\operatorname{Var}[\psi(X_{i},A_{i},R_{i},\theta_{0})]<\infty$. The results in this section derive necessary and sufficient conditions for $\tilde{\theta}_{n}$ to be regular, and further demonstrate that in order for $\tilde{\theta}_{n}$ to be regular and efficient, either $\psi=\psi^{\ast}$ or a fast-balancing condition involving $\psi(\cdot)$ needs to be satisfied. Therefore, finely stratified designs are not only sufficient to guarantee efficiency when estimating $\theta_{0}$ using the naïve method of moments estimator, but their fast-balancing property is also necessary.
@@
$\rho\bigg(\frac{1}{\sqrt{n}}\sum_{1\leq i\leq n}(A_{i}-\eta)\gamma(X_{i}),~N(0,V_{\gamma}^{\rm imb})\bigg|X^{(n)}\bigg)\xrightarrow{P}0$
for some deterministic variance $V_{\gamma}^{\rm imb}$.
@@
with all such assignments being drawn uniformly at random and independently across strata. It follows from Theorem 12.2.1 in lehmann2022testing combined with a subsequencing argument that Assumption 4.2 is satisfied with $V_{\gamma}^{\rm imb}=\eta(1-\eta)E[\operatorname{Var}[\gamma(X_{i})|S_{i}]]$.
@@
Revisiting Example 4.3, suppose treatment is assigned using the covariate adaptive randomization procedure described in ma2024new. Then it follows from Theorem 3.6 in their paper that, under appropriate assumptions,
@@
Therefore, by Markov’s inequality, Assumption 4.2 is satisfied with $V_{\gamma}^{\rm imb}=0$. By approximating square-integrable functions by square-integrable Lipschitz functions as in Lemma C.5 in cytrynbaum2023designing, it can be shown that the convergence holds for any square-integrable function $\gamma(\cdot)$.
@@
Revisiting Example 4.5, we note that, following Corollary 3.7 in cytrynbaum2024finely, we do not expect Assumption 4.2 to hold for re-randomization designs in general.
@@
which is indeed the case in finely stratified experiments when the treatment assignment mechanism satisfies Assumption 3.2.
@@ 5 Simulations
  3. Matched pairs, i.e., a finely stratified design with $(\ell,k)=(1,2)$, and the naïve method of moments estimator
  4. Matched tuples, i.e., a finely stratified design with $(\ell,k)=(2,4)$, and the naïve method of moments estimator
In Section 5.1, we present the model specifications and estimators for estimating the ATE as in Example 2.1. In Section 5.2, we present the model specifications and estimators for estimating the LATE as in Example 2.3. Section 5.3 reports the simulation results.
@@ 5.1 Average Treatment Effect
In this section, we present model specifications and estimators for estimating the ATE as in Example 2.1. Recall that in this case the moment function we consider is given by
@@
where $\mu_{0}(X_{i})=\sum_{1\leq l\leq 8}w_{l}\big(X_{i,l}+\frac{1}{3}(X_{i,l}^{2}-1)\big)$ for $w=(2,1,1,0.5,0.01,0.001,0.0001,\allowbreak 0.00001)$, $\mu_{1}(X_{i})=0.2+\mu_{0}(X_{i})$, $\epsilon_{i}\sim N(0,4)$, $(X_{i},\epsilon_{i})$, $1\leq i\leq n$ are i.i.d., and for each $1\leq i\leq n$, $(X_{i},\epsilon_{i})$ are independent. In each of the simulations to follow, when forming pairs/tuples or performing regression adjustment, we use only a subvector of the covariates $X_{i}$ consisting of the first $T$ covariates, for $T\in\{2,4,8\}$.
@@
The first estimator $\hat{\theta}_{n}^{\rm unadj}$ is the naïve method of moments estimator given by the solution to (6). The second and third estimators $\hat{\theta}_{n}^{\rm adj,1}$ and $\hat{\theta}_{n}^{\rm adj,2}$ are covariate-adjusted estimators which can be obtained as two-step method of moments estimators from solving the “augmented” moment equation (8) described in the discussion at the end of Section 2. $\hat{\theta}_{n}^{\rm adj,1}$ and $\hat{\theta}_{n}^{\rm adj,2}$ differ in the choice of basis functions used in the construction of the estimators $\hat{\mu}_{a}(x)$. Note that by the double-robustness property of the augmented estimating equation (8), it can be shown that the adjusted estimators $\hat{\theta}_{n}^{\rm adj,1}$, $\hat{\theta}_{n}^{\rm adj,2}$ are consistent and asymptotically normal regardless of the choice of estimators $\hat{\mu}_{a}(x)$, but consistency of $\hat{\mu}_{a}(x)$ to $\mu_{a}(x)$ would ensure that $\hat{\theta}_{n}^{\rm adj,1}$, $\hat{\theta}_{n}^{\rm adj,2}$ are efficient under i.i.d. assignment (robins1995analysis; tsiatis2008covariate; chernozhukov2017doubledebiasedneyman).
5.2 Local Average Treatment Effect
[…]
We consider the following three estimators for the LATE:
  • Unadjusted Estimator:
  • Adjusted Estimator 1:
Similarly to Section 5.1, $\hat{\theta}_{n}^{\rm unadj}$ solves (6) for the moment condition given in (5). The second and third estimators are covariate adjusted estimators which can be obtained as two-step method of moments estimators from solving an “augmented” version of the moment condition (5) (see, for instance, chernozhukov2018doubledebiased; jiang2022improving).
@@ 5.3 Simulation Results
Table 2 displays the ratio of the empirical MSE for each design/estimator pair relative to the MSE of the unadjusted estimator under i.i.d. assignment, also computed across $4000$ Monte Carlo replications. As expected given our theoretical results, we find that the empirical MSEs of the naïve unadjusted estimator under a matched pairs/tuples design closely match the empirical MSEs of the covariate adjusted estimators under i.i.d. assignment; this feature is particularly noteworthy given that the adjusted estimators are in fact correctly specified, and the correct specification would be unknowable in practice. We note that the MSE improvement of the unadjusted estimator with a matched pairs/tuples design relative to i.i.d. assignment is typically worse with 2 or 8 covariates than with 4 covariates: this phenomenon stems from the fact that the first 4 covariates are much stronger predictors of the control outcome than the last 4 covariates, which are almost uninformative. Although we have found in prior work (bai2024inference) that the matched pairs design delivers a lower MSE than the matched tuples design when the potential outcomes depend on the covariates linearly, we do not find that this is the case here with a nonlinear model. However, we do consistently find that the MSE with 8 covariates is smaller for the matched pairs design than the matched tuples design. This comparison illustrates that, although as discussed in Remark 3.1, our theoretical results imply matched pairs and matched tuples designs are not distinguishable asymptotically, their finite-sample properties may differ, especially when the number of covariates is large. In particular, note Assumption 3.2 is more stringent for matched tuples, for which $k=4$, than matched pairs, for which $k=2$.
Table 3 reports the coverage and average length of confidence intervals constructed using the estimator described in Section 3.1, again computed across $4000$ Monte Carlo replications. We find that the confidence intervals have appropriate coverage at all sample sizes and for any number of covariates. The average length of the intervals follows a pattern similar to what was observed for the MSE, again because the first 4 covariates are much stronger predictors of the control outcome than the last 4 covariates, which are almost uninformative. Although we found in Table 2 that the MSE of the matched tuples design was worse than that of matched pairs with 8 covariates, this does not seem to translate to longer confidence intervals for matched tuples: this may be due to the fact that the matched tuples variance estimator is a “within block” estimator, as described in Section 3.1. Similar findings have been reported by bai2024inference for related problems in factorial designs.
i.i.d. assignment Matched pairs Matched tuples
$n=100$
ATE 2 0.0083 0.0065 0.0051 0.0212 0.0068
4 0.0083 0.0090 0.0090 0.0073 0.0124
8 0.0083 0.0134 0.0131 0.0010 0.0209
LATE 2 0.0905 0.0436 0.0396 0.0341 0.0076
4 0.0905 0.0276 0.0264 0.0301 0.0427
8 0.0905 0.0401 0.0376 0.0264 0.0297
$n=200$
ATE 2 0.0075 0.0068 0.0074 0.0084 0.0081
4 0.0075 0.0052 0.0067 0.0040 0.0040
8 0.0075 0.0045 0.0055 0.0145 0.0051
LATE 2 0.0545 0.0326 0.0339 0.0361 0.0328
4 0.0545 0.0182 0.0217 0.0123 0.0144
8 0.0545 0.0159 0.0180 0.0447 0.0216
$n=400$
ATE 2 0.0001 0.0011 0.0009 0.0041 0.0055
4 0.0001 0.0001 0.0001 0.0001 0.0016
8 0.0001 0.0006 0.0006 0.0056 0.0041
LATE 2 0.0142 0.0064 0.0059 0.0179 0.0198
4 0.0142 0.0012 0.0015 0.0004 0.0020
8 0.0142 0.0021 0.0019 0.0129 0.0064
$n=1000$
ATE 2 0.0021 0.0037 0.0039 0.0014 0.0025
4 0.0021 0.0021 0.0022 0.0020 0.0006
8 0.0021 0.0021 0.0023 0.0023 0.0027
LATE 2 0.0118 0.0120 0.0123 0.0019 0.0083
4 0.0118 0.0055 0.0057 0.0051 0.0007
8 0.0118 0.0054 0.0058 0.0057 0.0059
$n=2000$
ATE 2 0.0013 0.0010 0.0010 0.0011 0.0025
4 0.0013 0.0015 0.0015 0.0003 0.0021
8 0.0013 0.0015 0.0015 0.0001 0.0014
LATE 2 0.0062 0.0035 0.0035 0.0018 0.0050
4 0.0062 0.0037 0.0038 0.0011 0.0055
8 0.0062 0.0038 0.0037 0.0008 0.0032
@@
Matched pairs Matched tuples
# of covariates ($T$) Coverage CI length Coverage CI length
$n=100$
ATE 2 0.9403 2.3191 0.9490 2.3224
4 0.9465 2.2744 0.9485 2.2830
8 0.9487 2.3532 0.9395 2.3672
LATE 2 0.9590 6.4075 0.9590 6.3865
4 0.9620 6.3184 0.9615 6.3176
8 0.9610 6.5497 0.9527 6.5311
$n=200$
ATE 2 0.9495 1.6401 0.9497 1.6418
4 0.9485 1.5990 0.9440 1.6049
8 0.9503 1.6529 0.9425 1.6617
LATE 2 0.9550 4.4046 0.9563 4.4264
4 0.9557 4.3315 0.9530 4.3381
8 0.9597 4.4833 0.9503 4.5038
$n=400$
ATE 2 0.9487 1.1602 0.9483 1.1614
4 0.9497 1.1255 0.9455 1.1285
8 0.9495 1.1609 0.9445 1.1667
LATE 2 0.9530 3.0955 0.9493 3.0963
4 0.9547 3.0122 0.9517 3.0110
8 0.9555 3.1284 0.9505 3.1239
$n=1000$
ATE 2 0.9485 0.7349 0.9550 0.7352
4 0.9500 0.7086 0.9505 0.7102
8 0.9417 0.7285 0.9430 0.7318
LATE 2 0.9507 1.9468 0.9560 1.9452
4 0.9515 1.8836 0.9507 1.8857
8 0.9467 1.9469 0.9463 1.9480
$n=2000$
ATE 2 0.9490 0.5194 0.9530 0.5193
4 0.9483 0.4998 0.9485 0.5004
8 0.9505 0.5121 0.9483 0.5139
LATE 2 0.9495 1.3714 0.9530 1.3721
4 0.9493 1.3250 0.9483 1.3258
8 0.9533 1.3663 0.9505 1.3658

### v8

@@ 4.1 Efficiency Bound
Given Theorem 4.1 we call $\mathbb{V}_{\ast}=\operatorname{Var}[\psi^{\ast}(X_{i},A_{i},R_{i},\theta_{0})]$ the efficiency bound for $\theta_{0}$, since our result shows that this is the lowest asymptotic variance attainable by any regular estimator under our assumptions. Indeed, it follows from Anderson’s lemma (Lemma 8.5 in van der Vaart, 1998) that the asymptotic loss of any regular estimator is bounded below by the loss under $N(0,\mathbb{V}_{\ast})$ for any “bowl-shaped” loss function (including, in particular, square loss). We note that our assumptions on the assignment mechanism preclude us from immediately appealing to standard semi-parametric convolution theorems (see, for instance, Theorem 25.20 in van der Vaart, 1989). Instead, we proceed by justifying an application of Theorem 3.1 in Armstrong (2022) combined with the convolution Theorem 3.11.2 in van der Vaart and Wellner (1996) to each $d_{\theta}$-dimensional parametric submodel separately, and then arguing that the supremum over all such submodels is attained by $\operatorname{Var}[\psi^{\ast}]$. A key observation is that in order to apply Theorem 3.1 in Armstrong (2022) to argue that the likelihood ratio process is locally asymptotically normal, the conditional information needs to settle down in the limit, which is guaranteed as long as Assumption 4.1 is satisfied.
@@
Following similar arguments as those in Remark 3.2, we can deduce that our efficiency bound agrees with well-known bounds for common parameters (like those presented in Examples 2.12.3) in the setting of i.i.d. assignment. For example, we have noted in the case of the ATE (Example 2.1) that (17) matches the efficiency bound under i.i.d. assignment derived in Hahn (1998). See Rafi (2023) and Armstrong (2022) for related results in the context of stratified and response-adaptive experiments. Straightforward calculation also implies that, for the quantile treatment effect (Example 2.2), the efficiency bound is given by
@@
which matches the efficiency bound under i.i.d. assignment derived in Firpo (2007) when the propensity score is set to $\eta$.
@@
The efficiency bound in Theorem 4.1 is attained by highly stratified experiments as in Theorem 3.1 if no additional covariates are available for estimation beyond the set of covariates $X_{i}$ used in the design. In practice, researchers may consider adjusting for additional baseline covariates in order to improve efficiency. Suppose additional covariates $W^{(n)}$ are available and Assumption 3.1 is modified such that
@@
Here, we comment on how Theorem 4.1 relates to prior efficiency bounds in experiments with general assignment mechanisms. For the case of estimating the ATE, Armstrong (2022) derives an efficiency bound over a very large class of assignment mechanisms, including even response-adaptive designs, and shows that the bound is attained when units are assigned to treatment (control) with conditional probability proportional to the conditional standard deviation of the potential outcome under treatment (control). This type of assignment is sometimes referred to as the Neyman allocation. On the other hand, our results show that this bound may be quite loose whenever the assignment proportions are restricted to be anything not equal to the Neyman allocation, which is, of course, unknown. For example, the bound is not informative about what can be achieved if the assignment proportions were set to one half regardless of whether or not the conditional outcome variances across treatment and control are equal. Such settings frequently arise in practice due to logistical constraints or the absence of pilot data with which to estimate conditional variances of potential outcomes under treatment and control. Furthermore, as argued in Cai and Rafi (2022), even if pilot data is available, these quantities may be estimated so poorly that exogenously constraining the assignment proportions to one half leads to more efficient estimates of the ATE in practice. Motivated by such concerns, Rafi (2023) derives an efficiency bound for the ATE over the class of “coarsely-stratified” assignment mechanisms studied in Bugni et al. (2019), where the stratum-level assignment proportions are restricted a priori by the experimenter. This framework, however, rules out highly stratified designs. Finally, we once again emphasize that our analysis, unlike these other papers, applies to a general class of treatment effect parameters, including the ATE as a special case.
@@
If we additionally impose that $\eta(X_{i})$ takes on a finite set of values $\{\eta_{1},\dots,\eta_{S}\}$, then this bound could be achieved by separately implementing a highly stratified experiment over each set $\{i:\eta(X_{i})=\eta_{s}\}$ for $1\leq s\leq S$. In other words, separately within each stratum defined by the units for which $\eta(X_{i})=\eta_{s}$, employ the assignment mechanism described in Assumptions 3.13.2 with $\ell/k=\eta_{s}$. For more general functions $\eta(\cdot)$, we conjecture one could employ the local randomization procedure in Cytrynbaum (2023b).
@@ 4.2 The Necessity of “Fast-Balancing”
In this subsection, we provide conditions under which the fast-balancing condition described in (14) is a necessary condition for efficient estimation of $\theta_{0}$ “by design.” As a supplement, we also provide necessary and sufficient conditions for an asymptotically linear estimator to be regular (in the sense of (S.16) in Supplement A.2) for a large class of treatment assignment mechanisms. Concretely, given an assignment mechanism, suppose $\tilde{\theta}_{n}$ is an asymptotically linear estimator for $\theta_{0}$ in the sense that
@@
where $E[\psi(X_{i},A_{i},R_{i},\theta_{0})]=0$ and $\operatorname{Var}[\psi(X_{i},A_{i},R_{i},\theta_{0})]<\infty$. The results in this section derive necessary and sufficient conditions for $\tilde{\theta}_{n}$ to be regular, and further demonstrate that in order for $\tilde{\theta}_{n}$ to be regular and efficient, either $\psi=\psi^{\ast}$ or a fast-balancing condition involving $\psi(\cdot)$ needs to be satisfied. Therefore, highly stratified designs are not only sufficient to guarantee efficiency when estimating $\theta_{0}$ using the naïve method of moments estimator, but their fast-balancing property is also necessary.
@@
$\rho\bigg(\mathcal{L}\Big(\frac{1}{\sqrt{n}}\sum_{1\leq i\leq n}(A_{i}-\eta)\gamma(X_{i})\Big|X^{(n)}\Big),~N(0,V_{\gamma}^{\rm imb})\bigg)\xrightarrow{P}0$
for some deterministic variance $V_{\gamma}^{\rm imb}$, where $\mathcal{L}(\cdot|X^{(n)})$ denotes the conditional distribution given $X^{(n)}$.
@@
with all such assignments being drawn uniformly at random and independently across strata. It follows from Theorem 12.2.1 in Lehmann and Romano (2022) combined with a subsequencing argument that Assumption 4.2 is satisfied with $V_{\gamma}^{\rm imb}=\eta(1-\eta)E[\operatorname{Var}[\gamma(X_{i})|S_{i}]]$.
@@
Revisiting Example 4.3, suppose treatment is assigned using the covariate adaptive randomization procedure described in Ma et al. (2024). Then it follows from Theorem 3.6 in their paper that, under appropriate assumptions,
@@
Therefore, by Markov’s inequality, Assumption 4.2 is satisfied with $V_{\gamma}^{\rm imb}=0$. By approximating square-integrable functions by square-integrable Lipschitz functions as in Lemma C.5 in Cytrynbaum (2023b), it can be shown that the convergence holds for any square-integrable function $\gamma(\cdot)$.
@@
Revisiting Example 4.5, we note that, following Corollary 3.7 in Cytrynbaum (2024), we do not expect Assumption 4.2 to hold for re-randomization designs in general.
@@
which is indeed the case in highly stratified experiments when the treatment assignment mechanism satisfies Assumption 3.2.
@@ 5 Simulations
  3. Matched pairs, i.e., a highly stratified design with $(\ell,k)=(1,2)$, and the naïve method of moments estimator
  4. Matched quadruplets, i.e., a highly stratified design with $(\ell,k)=(2,4)$, and the naïve method of moments estimator
In Section 5.1, we present the model specifications and estimators for estimating the ATE as in Example 2.1. Supplement B contains the model specifications and estimators for estimating the LATE as in Example 2.3. Section 5.2 reports the simulation results for the MSE. The tables and discussion for bias and coverage are contained in Supplement B.
@@ 5.1 Average Treatment Effect
In this section, we present model specifications and estimators for estimating the ATE as in Example 2.1. The Recall that in this case the moment function we consider is given by
@@
where $\mu_{0}(X_{i})=\sum_{1\leq l\leq 8}w_{l}\big(X_{i,l}+\frac{1}{3}(X_{i,l}^{2}-1)\big)$ for $w=(2,1,1,0.5,0.01,0.001,0.0001,\allowbreak 0.00001)$, $\mu_{1}(X_{i})=0.2+\mu_{0}(X_{i})$, $\epsilon_{i}\sim N(0,4)$, $(X_{i},\epsilon_{i})$, $1\leq i\leq n$ are i.i.d., and for each $1\leq i\leq n$, $(X_{i},\epsilon_{i})$ are independent. In each of the simulations to follow, when forming pairs/quadruplets or performing regression adjustment, we use only a subvector of the covariates $X_{i}$ consisting of the first $T$ covariates, for $T\in\{2,4,8\}$.
@@
The first estimator $\hat{\theta}_{n}^{\rm unadj}$ is the naïve method of moments estimator given by the solution to (6). The second and third estimators $\hat{\theta}_{n}^{\rm adj,1}$ and $\hat{\theta}_{n}^{\rm adj,2}$ are covariate-adjusted estimators which can be obtained as two-step method of moments estimators from solving the “augmented” moment equation (8) described in the discussion at the end of Section 2. $\hat{\theta}_{n}^{\rm adj,1}$ and $\hat{\theta}_{n}^{\rm adj,2}$ differ in the choice of basis functions used in the construction of the estimators $\hat{\mu}_{a}(x)$. Note that by the double-robustness property of the augmented estimating equation (8), it can be shown that the adjusted estimators $\hat{\theta}_{n}^{\rm adj,1}$, $\hat{\theta}_{n}^{\rm adj,2}$ are consistent and asymptotically normal regardless of the choice of estimators $\hat{\mu}_{a}(x)$, but consistency of $\hat{\mu}_{a}(x)$ to $\mu_{a}(x)$ would ensure that $\hat{\theta}_{n}^{\rm adj,1}$, $\hat{\theta}_{n}^{\rm adj,2}$ are efficient under i.i.d. assignment (Robins et al., 1995; Tsiatis et al., 2008; Chernozhukov et al., 2017).
@@ 5.2 Simulation Results
We focus on the MSE and leave the discussion for bias and coverage to Supplement B. Table 1 displays the ratio of the empirical MSE for each design/estimator pair relative to the MSE of the unadjusted estimator under i.i.d. assignment, computed across $4000$ Monte Carlo replications. As expected given our theoretical results, we find that the empirical MSEs of the naïve unadjusted estimator under a matched pairs/quads design closely match the empirical MSEs of the covariate adjusted estimators under i.i.d. assignment; this feature is particularly noteworthy given that the adjusted estimators are in fact correctly specified, and the correct specification would be unknowable in practice. We note that the MSE improvement of the unadjusted estimator with a matched pairs/quads design relative to i.i.d. assignment is typically worse with 2 or 8 covariates than with 4 covariates: this phenomenon stems from the fact that the first 4 covariates are much stronger predictors of the control outcome than the last 4 covariates, which are almost uninformative. Although we have found in prior work (Bai et al., 2024c) that the matched pairs design delivers a lower MSE than the matched quads design when the potential outcomes depend on the covariates linearly, we do not find that this is the case here with a nonlinear model. However, we do consistently find that the MSE with 8 covariates is smaller for the matched pairs design than the matched quads design. This comparison illustrates that, although as discussed in Remark 3.1, our theoretical results imply matched pairs and matched quads designs are not distinguishable asymptotically, their finite-sample properties may differ, especially when the number of covariates is large. In particular, note Assumption 3.2 is more stringent for matched quads, for which $k=4$, than matched pairs, for which $k=2$.